Grade 12th passPhysical Chemistry

Calculate the number of kJ of heat necessary to raise the temperatureof 60.0 g of aluminium from 35°C to 55°C. Molar heat capacity of Al is 24 J ..???????

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Given mass Al =60g

Molar mass Al = 27g/mol

so,mole=given mass /molar mass

60/27 =20/9

A/C question ,

Molar heat capacity (C) = 24J/mol/k

Temperature change (∆T) =(55° C 35° C)

=20°C or 20 k

Heat {Q} = n.c.∆T

=20/9 ×24×20J

=1066.67 J

=1.0667KJ


Thanks